OpenStudy (anonymous):

Check my answer?? solve for x 6+1/6x=11 help! fractions confuse me

OpenStudy (kewlgeek555):

Did you know that fractions are division signs? Yeah, they just invented fractions because imagine 1/6. That is a long decimal! I am sure decimals might be easier to operate with, though. Anywho, this is basically asking \[6+(1\div6)x=11\]Now you have to work backwards. Since the x is to right, lets start in the left.\[11-6=5\]\[(1\div6)x=5\]Now, since you cannot do 11 * 1 and then 11 * 6, just divide 1 and 6.\[0.166666667x = 5\] All you have to do is \[5\div0.166666667 = x\]You can do it on a calculator.

OpenStudy (anonymous):

so x=29.999999994? i don't think i did it right.

OpenStudy (doc.brown):

\[6+\frac{1}{6}x=11\]\[6-6+\frac{1}{6}x=11-6\]\[\frac{1}{6}x=5\]\[6\times\frac{1}{6}x=5\times6\]\[\frac{6}{6}x=30\]\[1x=30\]\[x=30\]

OpenStudy (anonymous):

oooh i get it now. thank you sooo much!

OpenStudy (kewlgeek555):

@sammysweets teachers I think do decimals rounded to the tenths or hundreth place, but since I used the exact decimal, I got the exact answer. But it is 30, rounded.

OpenStudy (doc.brown):

A fraction tells you how many equal parts you have, and how many of them make up a whole. \[\frac{1}{2}apple\]means you have one part of an apple, but you require two to make an entire apple. You can also view it as\[\frac{1}{2}\times apple\]or\[\frac{1\times apple}{2}\]or\[\frac{apple}{2}\] Cut an apple in two, what do you get?